Help me with this question

A spring is stretched by 60cm which transfers 18 J of energy to its elastic potential energy store. What is the spring constant?

Help me please URGENTLY
1
Expert's answer
2020-09-30T12:27:07-0400

Potential energy of spring is "U = \frac{k x^2}{2}", where "k" is spring constant, and "x" is distance stretched.

From the last formula, "k = \frac{2 U}{x^2} = \frac{2 \cdot 18 J}{(0.6m)^2} = 100 \frac{N}{m}".